OpenStudy (anonymous):

Write an expression for the slope of a line passing through the points (0,-3) and (x,-9). What values of x will produce a positive slope? What values of x will produce a negative slope? What value of x will produce a vertical line? Can a horizontal line be formed from these two points? Why or why not?

OpenStudy (amistre64):

x < 0 would make a positive slope; at x = 0 we have an undefined slope

OpenStudy (amistre64):

x = (-inf,0) produces a positive slope x = (0,inf) makes it a negative slope; and at x=0 we get a vertical line that is undefined

OpenStudy (anonymous):

So what will the expression be for the first part of the problem?

OpenStudy (amistre64):

what do you mean?

OpenStudy (anonymous):

Well, I'm thinking I have to put this into a formula like the "two point formula" since I was given two points. No?

OpenStudy (amistre64):

the slope formula for these to points as is: -9+3 ----- = m x - 0

OpenStudy (amistre64):

or simply: -6 --- = m x

OpenStudy (anonymous):

Or maybe this one: y=mx + b?

OpenStudy (amistre64):

y+3 -6 ---- = --- x-0 x y+3 = (-6/x)(x-0) y = -9 is what they tunrs out to, but I dont think that is right